OpenStudy (anonymous):

Find the cube roots of 27(cos 279° + i sin 279°).

jimthompson5910 (jim_thompson5910):

Use the formulas found on this page http://www.mathxtc.com/Downloads/NumberAlg/files/NthRootsComplex.pdf to get the following First cube root \[\Large \sqrt[3]{27}\left(\cos\left(\frac{279}{3}\right)+i\sin\left(\frac{279}{3}\right)\right)\] \[\Large 3\left(\cos(93)+i\sin(93)\right)\] So the first cube root is \[\Large 3\left(\cos(93)+i\sin(93)\right)\] ------------------------------------------------------------------------------ Second cube root \[\Large \sqrt[3]{27}\left(\cos\left(\frac{279}{3}+\frac{360}{3}\right)+i\sin\left(\frac{279}{3}+\frac{360}{3}\right)\right)\] \[\Large 3\left(\cos(93+120)+i\sin(93+120)\right)\] \[\Large 3\left(\cos(213)+i\sin(213)\right)\] The second cube root is \[\Large 3\left(\cos(213)+i\sin(213)\right)\] ------------------------------------------------------------------------------ Third cube root \[\Large \sqrt[3]{27}\left(\cos\left(\frac{279}{3}+\frac{2*360}{3}\right)+i\sin\left(\frac{279}{3}+\frac{360}{3}\right)\right)\] \[\Large 3\left(\cos(93+240)+i\sin(93+240)\right)\] \[\Large 3\left(\cos(333)+i\sin(333)\right)\] and the third cube root is \[\Large 3\left(\cos(333)+i\sin(333)\right)\] ------------------------------------------------------------------------------